Jacki is seven years older than Penny. Twelve years ago, Penny was half as old as Jacki is now. How old is Jacki now?
Answer by VFBundy(438)

Jacki's age now = x
Penny's age now = x - 7

2((x - 7) - 12) = x

2(x - 19) = x

2x - 38 = x

x - 38 = 0

x = 38

Jacki's age now = x = 38

Check it:

If Jacki is 38 now, Penny is 31 now. (Because Jacki is seven years older than Penny.)

"Twelve years ago, Penny was half as old as Jacki is now." Twelve years ago, Penny was 19. (Because Penny is 31 now.) Jacki is 38 now. Nineteen is half of 38, so that checks.